Olivier Schiffmann: The Hall algebra of an elliptic curve

We describe the Hall algebra of an elliptic curve by generators and relations, and show that the group of autoequivalences of the derived category acts on its Drinfeld double. We also give a precise link with the theory of symmetric functions. This is joint work with I. Burban.